Terri drove for 3 hours before lunch. After lunch she drove 4 more hours and averaged 15mph more than before lunch. If her total distance was 410 miles, what was her average speed before lunch?
Answer by ikleyn(52786)

.

Write the total distance equation

    3B + 4*(B+15) = 410  miles  total.

where B is her average rate before lunch, in miles per hour.


Simplify the equation and find B

    3B + 4B + 60 = 410

      7B         = 410 - 60

      7B         =   350

       B         =   350/7 = 50.


ANSWER.  Terri's average rate before lunch was 50 miles per hour.

Solved.